OpenStudy (anonymous):

if concrete weighs 160lbs per cubic ft, what is the weight of a concrete circular slab 4in thick and 15 ft across

OpenStudy (anonymous):

Cubic ft l * w * h 160 = l * w * h 160 = 15 * w * 1/3 you have to put all of these in ft. 160 = 5w w = 32 ft. I assume you meant width and not weight.

OpenStudy (anonymous):

its weight

OpenStudy (anonymous):

OOPS sorry.. didn't read circular disregard what I wrote above.

OpenStudy (anonymous):

cylinder = pi r^2 h cylinder = 3.14 * (7.5)^2 * 1/3 cylinder = 58.875 ft^3 So it weighs 160 lbs for each ft^3 of concrete. You have 58.875 ft^3 so 160* 58.875 = 942 pounds.